OpenStudy (laddiusmaximus):

cosx/1-sinx -1/cosx = tanx what happens to the 1-sinx? I know 1/cosx becomes sec, but what then?

OpenStudy (amistre64):

it should "post" when you hit "post" .....

OpenStudy (anonymous):

flip the order and you will have : -1-sinx/cosx = -1/cosx -sinx/cosx sinx/cosx = tanx so; -1/cosx - tanx = -secx-tanx

OpenStudy (anonymous):

I guess

OpenStudy (amistre64):

cos - 1 sin ----- --- = --- is this the equation? 1-sin cos cos

OpenStudy (laddiusmaximus):

yes

OpenStudy (laddiusmaximus):

would algebra be the first step? common denominators?

OpenStudy (amistre64):

give it a shot :)

OpenStudy (laddiusmaximus):

I thought at first using the reciprocal identity for the 1/cos

OpenStudy (laddiusmaximus):

okay now I have sin^2x/1-sinxcosx

OpenStudy (laddiusmaximus):

which then becomes sinx/1-cosx

OpenStudy (amistre64):

thinking......(12% loaded)

OpenStudy (amistre64):

thinking done... got it :)

OpenStudy (amistre64):

get like denoms is good; multiply cos/(1-sin) by (cos/cos) to get what?

OpenStudy (amistre64):

and 1/cos get multiplied by (1-sin)/(1-sin) to get??

OpenStudy (amistre64):

cos^2 - (1-sin) ------------ right? cos(1-sin)

OpenStudy (amistre64):

work the top out to: cos^2 + sin -1 ; cos^2 = 1-sin^2 1-sin^2 + sin - 1 ; 1-1 = 0 sin - sin^2 ; factor out a "sin" sin(1-sin) is the top; cancel out (1-sin) top to bottom to get: sin/cos = tan....TADA!!!

OpenStudy (amistre64):

you follow that alright?

OpenStudy (laddiusmaximus):

yeah but I ended up with cos^2x-1/1-sinxcosx

OpenStudy (laddiusmaximus):

then sin^2x/1-sinxcosx

OpenStudy (amistre64):

the bottom aint 1-sincos; its cos(1-sin)

OpenStudy (amistre64):

cos cos - (1-sin) -------- ------- right?? cos(1-sin) cos(1-sin)

OpenStudy (amistre64):

to get like denoms; you need to multiply each side by a convenient form of (1)

OpenStudy (amistre64):

cos/cos = 1 AND (1-sin)/(1-sin) = 1

OpenStudy (laddiusmaximus):

ohhhh duh

OpenStudy (amistre64):

lol ..... mathmatikers are tricky like that :)

OpenStudy (laddiusmaximus):

then I start canceling out like numbers right?

OpenStudy (amistre64):

there some manipulations first.

OpenStudy (amistre64):

the bottom is good; so play with the top.

OpenStudy (laddiusmaximus):

so cosx cosx - (1-sinx) becomes cos^2x -(1-sinx)

OpenStudy (amistre64):

good; now what multiply out the -(1-sin). what do we get?

OpenStudy (laddiusmaximus):

1+sinx

OpenStudy (amistre64):

good; and how can we play with cos^2 to make it in terms of "sin"? what does cos^2 =?

OpenStudy (laddiusmaximus):

sin^2x

OpenStudy (laddiusmaximus):

pythagorean identity?

OpenStudy (amistre64):

close.... yes, pyth identity. sin^2 + cos^2 = 1 solve for cos^2, what do we get? subtract sin^2 from each side: cos^2 = 1-sin^2

OpenStudy (amistre64):

our top becomes what with this information?

OpenStudy (laddiusmaximus):

1-sin^2x-1+sinx

OpenStudy (amistre64):

not ..... -1+sin. remember we already took care of that part. It was equal to: +1+sin do you see that? do you understand how we did it?

OpenStudy (laddiusmaximus):

but I thought -(1-sinx) would become -1+sinx - times one is -1 and - times -sinx would become +sinx

OpenStudy (anonymous):

lol. now I'm lost ._.

OpenStudy (amistre64):

you know what.... YOU are correct. I was wrong.... carry on please :)

OpenStudy (anonymous):

who?

OpenStudy (amistre64):

Laddi was right. And im glad Laddi caught it; have confidence in yourself :)

OpenStudy (anonymous):

sorry for the wrong answer laddi!

OpenStudy (amistre64):

1-sin^2-1+sin what can we do with this?

OpenStudy (anonymous):

replace 1-sin^2 with cose^2

OpenStudy (amistre64):

sst; only if we wanna work backwards... lets try to work forwards :)

OpenStudy (anonymous):

but why?

OpenStudy (laddiusmaximus):

does it become sinx?

OpenStudy (amistre64):

lets rewrite it this way: 1-1+sin-sin^2 now what do we do?

OpenStudy (laddiusmaximus):

well 1-1 becomes 0, so sin-sin^2 is left.

OpenStudy (amistre64):

very good; how can we work this? can we factor anything out to make it look more like the bottom of the fraction? remember the bottom looks like this: cos(1-sin)

OpenStudy (amistre64):

think of the top as: x-x^2 if it helps

OpenStudy (laddiusmaximus):

sinx can factor out 1-sinx

OpenStudy (anonymous):

why don't you just flip cos/1-sin so it will be -(1-sin)/cos?

OpenStudy (amistre64):

very good :) that leaves us with: sin(1-sin) sin -------- = --- cos(1-sin) cos can you see that we can do now?

OpenStudy (amistre64):

sst; because you change its value if you do that and make a different equation. does 2/x = x/2 for all "x"?

OpenStudy (anonymous):

no ofcourse not it's equal to -x/2 ^^

OpenStudy (anonymous):

wait, I get what you mean now ^^"

OpenStudy (laddiusmaximus):

wait a minute how did (1-sin) get up top? I had sinx+sin^2x

OpenStudy (amistre64):

lol .... at least one of us does :)

OpenStudy (anonymous):

see, you flipped it!

OpenStudy (anonymous):

I knew it

OpenStudy (amistre64):

we had: 1-1+sin-sin^2 = sin-sin^2 = sin(1-sin)

myininaya (myininaya):

find common denominator which you guys did! Good! [ (cosx)^2-1]/[(1-sin(x))(cosx)]

myininaya (myininaya):

Now 1=(cosx)^2+(sinx)^2

myininaya (myininaya):

Replace 1 with that

myininaya (myininaya):

The (cosx)^2 cancel out

OpenStudy (anonymous):

there's still more? I thought the question was answered

myininaya (myininaya):

The nemerator has -(sinx)^2+sinx

OpenStudy (laddiusmaximus):

ok confusion. 1-sin^2x-1+sinx becomes what?

OpenStudy (amistre64):

myin; your equation is skewed look at what you posted again :)

myininaya (myininaya):

You can factor out a (sinx) leaving with sinx(-sinx+1)

OpenStudy (laddiusmaximus):

sin^2x+sinx

OpenStudy (amistre64):

Laddi: (-sin^2) + sin dont forget you (-)

myininaya (myininaya):

You can do a cancellation leaving you with the desired reult

OpenStudy (amistre64):

1 (-)sin^2 (-)1 (+)sin becomes: (-)sin^2 (+)sin

OpenStudy (laddiusmaximus):

okay, then what?

myininaya (myininaya):

Factor out what they have in common which is a sinx

OpenStudy (amistre64):

then we factor out a (+sin) to get: sin(1-sin) correct?

OpenStudy (laddiusmaximus):

okay so -sinx is left on top?

myininaya (myininaya):

The numerator gives you sinx(-sinx+1)

OpenStudy (anonymous):

no sinx(sinx - 1)

OpenStudy (laddiusmaximus):

wait a minute. okay right now I have -sin^2x +sinx/cosx(1-sinx)

myininaya (myininaya):

No ss the sinx is negative and the 1 is postive

OpenStudy (amistre64):

good Laddi;

OpenStudy (laddiusmaximus):

so -sin^2x factors out the +sinx?

OpenStudy (amistre64):

yes...

OpenStudy (laddiusmaximus):

so im left with -sinx/cosx(1-sinx)

myininaya (myininaya):

You should be left with [(sinx(-sinx+1)]/[cosx(1-sinx)]

myininaya (myininaya):

Does anything cancel?

OpenStudy (amistre64):

(-sin^2 +sin)/cos(1-sin) becomes... sin(1-sin) / cos(1-sin)

OpenStudy (anonymous):

I'll just wacth ^-^

myininaya (myininaya):

-sinx+1=1-sinx So that part in you numerator cancels out with that part in the denominator

OpenStudy (amistre64):

Laddi: Does -sin^2 + sin = sin(1-sin) ???

OpenStudy (anonymous):

-sin(sin -1)

OpenStudy (laddiusmaximus):

I didnt think it did.

OpenStudy (anonymous):

or sin(-sin + 1) = sin (1-sin)